# MetaFood3D: Large 3D Food Object Dataset with Nutrition Values

### v1

@@ 4.4 Food Portion Estimation
Baseline 151.85 845.69 221.37 1287.25
Stereo Reconstruction [9] 135.96 210.90 271.78 244.55
Voxel Reconstruction [11] 123.34 104.07 190.38 145.31
RGB Only (ResNet50) [68] - - 1932.01 1124.9
Density Map Only (ResNet50) [79] - - 1100.39 663.43
Density Map Summing [46] - - 436.12 142.44
3D Assisted Portion Estimation [78] 195.92 79.33 260.79 102.25
Table 5: Comparison of image-based dietary assessment methods on the MetaFood3D dataset.
The results presented in Table 5 demonstrate the performance of different classes of existing methods on our MetaFood3D dataset. The 3D Assisted Portion Estimation method [78] achieves the lowest V-MAPE and E-MAPE. However, it is important to note that this method has higher V-MAE and E-MAE values compared to some other approaches, which suggests that the incorporation of 3D object information can lead to improved performance in terms of percentage errors, but further research may be needed to reduce the absolute errors. The MetaFood3D dataset provides a valuable resource for developing and evaluating various dietary assessment techniques, and the presented results highlight the potential for further improvements in food portion estimation accuracy.
@@ 5 Conclusion
In this paper, we present MetaFood3D, a food-specific 3D object dataset to advance dietary analysis. This new dataset provides a robust benchmark for developing and evaluating 3D vision algorithms for real-world scenarios. The dataset features diverse intra-class variations, detailed nutrition annotations and rich multimodal data. Experimental results demonstrate great potential for using our dataset in various downstream tasks related to food image analysis.

### v2

@@ 4.4 Food Portion Estimation
Baseline 165.75 836.50 214.55 1135.93
Stereo Reconstruction [9] 153.58 214.95 262.07 244.80
Voxel Reconstruction [11] 120.16 96.31 174.45 130.16
RGB Only [70] - - 1500.23 370.9
Density Map Only [83] - - 1098.87 654.33
Density Map Summing [46] - - 426.68 146.18
3D Assisted Portion [84] 186.45 83.26 287.11 132.42
MPF3D [47] 62.60 41.43 77.98 68.05
Table 5: Comparison of image-based dietary assessment methods on the MetaFood3D dataset. The last couple of rows are methods that utilize the 3D models in the MetaFood3D dataset for portion estimation
The results presented in Table 5 highlight the performance of different classes of existing methods on our MetaFood3D dataset. The MPF3D [47] demonstrates the importance of 3D information for portion estimation outperforming stereo-based, depth-based, and network-based methods on all metrics. The 3D Assisted Portion Estimation method [84] achieves the second lowest V-MAPE and E-MAPE. The performance improvement offered by the 2 methods of portion estimation that utilize 3D information from our dataset underscores the important role that 3D food models play in the field of food portion estimation. Thus, the MetaFood3D dataset provides a valuable resource for developing and evaluating various dietary assessment techniques.
@@ 5 Conclusion
In this paper, we present MetaFood3D, a food-specific 3D object dataset to advance food computing and 3D computer vision. This new dataset provides a robust benchmark for developing and evaluating 3D vision algorithms for real-world scenarios. The dataset features diverse intra-class variations, detailed nutrition annotations and rich multimodal data. Experimental results demonstrate the strong capabilities of our dataset in food portion estimation, synthetic eating occasion simulation, and 3D food object generation.